how do I solve this rational expression
9 / x+5 + x / x-3 = 5
my answer is x = 2 and x = 7/3

Answer by stanbon(75887)

solve this rational expression
9 / (x+5) + x / (x-3) = 5
--------------------
Multiply thru by (x+5)(x-3) = x^2+2x-15 to get:
---------------------------------------
9(x-3) + x(x+5) = 5(x^2+2x-15)
9x - 27 + x^2 + 5x = 5x^2 + 10x - 75
---------------------------------------
4x^2 -4x -48 = 0
x^2 - x - 12 = 0
(x-4)(x+3) = 0
x = 4 or x = -3
=====================
Cheers,
Stan H.